Prove that the sum of two convergent sequence is convergent.

1
Expert's answer
2021-11-16T11:21:57-0500

Let A and B be the points of convergence of the two respective series. The convergence of the two series implies that given any "\varepsilon>0" , there exists an integer "N_{0}" such that for any "N \geq N_{0}" , we have

"\begin{aligned}

&\left|A-\sum_{n=1}^{N} a_{n}\right|<\varepsilon \\

&\left|B-\sum_{n=1}^{N} b_{n}\right|<\varepsilon

\end{aligned}"

We claim that "\sum_{n=1}^{\infty} \frac{a_{n}+b_{n}}{2}" converges to "\frac{A+B}{2}" . Indeed, for all "N \geq N_{0}" , we can use the triangle inequality to get

"\left|\frac{A+B}{2}-\sum_{n=1}^{N} \frac{a_{n}+b_{n}}{2}\right| \leq \frac{1}{2}\left|A-\sum_{n=1}^{N} a_{n}\right|+\frac{1}{2}\left|B-\sum_{n=1}^{N} b_{n}\right|<\varepsilon"

Hence, the sum of two convergent sequence is convergent.